You decide to sell a piece ofland that you inherited a few years ago. The buyer agrees to either pay you $24,000 at closing or according to the schedule below (payments are assumed to be at the beginning of each of the 5 years). You do not really need the money today and are planning either way to just let the funds accumulate in an account earning 7%. However, you do plan to use the funds at the end of 5 years after closing on the land transaction to build a new home. Which altemative provides the highest future value at the end of 5 years? Beginning of Year Cash Flow $2,000 $4,000 6,(M $8,000 $10,000 2
Answer #1
Future value = present value * FVIFA(Interest rate , n)
For 24000, the Future value = 24000 * FVIFA(7% , 5) = 24000 * 5.751 = $138024
For Yearly payment, the Future Value = 2000 * FVIFA(7%,4) + 4000 * FVIFA(7%,3) + 6000 * FVIFA(7%,2) + 8000 * FVIFA(7%,1) + 10000 =
2000 * 4.440 + 4000 * 3.215 + 6000 * 2.070 + 8000 * 1 + 10000 = $52160
Selected Alt. : Lumpsum payment at present of $24000 would provide highest future value at the end of 5th year.